OpenStudy (anonymous):

How many possible three-digit passwords can be formed using digits 0 through 9 if digits are repeated? 30 , 720 , 1,000

OpenStudy (amistre64):

10 options for the 1st number 10 options each for those 10; 10*10 and 10 options each for those: 10*10*10

OpenStudy (anonymous):

which is the same as asking "how many three digit numbers are there?"

OpenStudy (lollylau):

what do you mean by "digits are repeated"?

OpenStudy (amistre64):

000 to 999

OpenStudy (lollylau):

does that mean that at least two of them must be the same?

OpenStudy (anonymous):

idk its on my math test

OpenStudy (lollylau):

-.-

OpenStudy (lollylau):

go ask your teacher

OpenStudy (amistre64):

lol, i hate interpreting word problems

OpenStudy (amistre64):

do digits have to be repeated? or CAN they be repeated?

OpenStudy (anonymous):

no , she doesnt even know

OpenStudy (lollylau):

how can she don't know???????

OpenStudy (amistre64):

most likely I would say it means that the option is left open that they CAN be repeated, but is not a requirement

OpenStudy (anonymous):

bc shes dumb like that , its an online class & shes not even a math teacher

OpenStudy (lollylau):

but its given =.=

OpenStudy (amistre64):

lets not degrade the teachers please.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Here is a good way to think of such things, I find...Let some arbitrary password be \[x_1x_2x_3\] (where we mean digits next to each other, not multiplication). So we have \[x_1,x_2,x_3 \in \{0,1,2,...,9\}\]Now, each position has the choice of 10 digits, 0 through 9 inclusive. So the answer is 10x10x10 = 10^3 = 1000.

OpenStudy (anonymous):

If the password was n digits long, then the number of possible choices is \[10^n\]

OpenStudy (anonymous):

amistre64 has a good way of thinking about it, since your first digit has 10 options, and for each of those first digits there are 10 options for your second digit. So there are 10*10=100 options for two digits. Now for each of those 100 options, there are 10 options for your third digit. So for any three digits, there are 100*10 options.

OpenStudy (anonymous):

logically speaking... 3 digit password starts from 000 to 999.... now u can count how many of them are between 000 & 999 inclusive. It will include all possible repeating numbers.